Consider the distribution of Grades in Dr. Smith's class:
Gender
Grade Male Female
A 18 12
B 30 30
C 53 27
D 12 8
F 7 3
120 80
If I select a student at random, what is the probability that:
a. the student earned an A?
b. the student earned at least a C?
c. the student earned a D, given that it's a male student?
Not a textbook question, thank you for any assistance :)

Answer by stanbon(57984)

Consider the distribution of Grades in Dr. Smith's class:
Gender
Grade Male Female
A 18 12
B 30 30
C 53 27
D 12 8
F 7 3
120 80
If I select a student at random, what is the probability that:
a. P(the student earned an A) = 30/200 = 3/20
b. P(the student earned at least a C) = (80+60+30)/200 = 17/20
c. P(the student earned a D, given that it's a male student) = 12/120 = 1/10
================
Cheers,
Stan H.